find the length of the arc of the curve 9y2= (x2+2)3 from the point where x=0 To the point where x = 2.


1
Expert's answer
2022-06-07T08:46:14-0400
"3y=\pm(x^2+2)^{3/2}"

"y=\pm\dfrac{1}{3}(x^2+2)^{3/2}"

Take "y=\dfrac{1}{3}(x^2+2)^{3/2}"


"y'=\dfrac{1}{3}(\dfrac{3}{2})(x^2+2)^{1/2}(2x)=x\sqrt{x^2+2}"


"L=2\displaystyle\int_{0}^{2}\sqrt{1+(x\sqrt{x^2+2})^2}dx"

"=2\displaystyle\int_{0}^{2}\sqrt{1+2x^2+x^4}dx"

"=2\displaystyle\int_{0}^{2}\sqrt{(1+x^2)^2}dx"

"=2\displaystyle\int_{0}^{2}(1+x^2)dx"

"=2[x+\dfrac{x^3}{3}]\begin{matrix}
 2 \\
 0
\end{matrix}=2(2+\dfrac{8}{3}-0)=\dfrac{28}{3}"